patricia is twice as old as laura. in three years, the sum of their ages will be 52. how old is patricia?
Answer by addingup(3677)

L = 1/2P
P+3+L+3 = 52
P+3+1/2P+3 = 52
1.5P+6 = 52
1.5P = 46
P = 30.667 or 30 years and 6 months.